For the parabola passing through the points (1,1) (3,13) (5,49)
What is the equation in the form y=ax^2+bx+c?
What is the vertex of the parabola?

Answer by ewatrrr(24785)

 
Hi,
For the parabola passing through the points (1,1) (3,13) (5,49)
a + b + c = 1
9a + 3b + c = 13
25a + 5b + c = 49
a = 3, b = -6 and c = 4
y = 3x^2 -6x + 4
OR
y = 3(x-1)^3 + 1 Vertex(1,1)